OpenStudy (anonymous):

Factoring trinomial completly.... y^3-12y^2+32y

OpenStudy (anonymous):

y(y-4)(y-8)

OpenStudy (anikhalder):

Simplifying the exp : =y(y^2 -12y +32) =y(y^2 -4y -8y +32) =y{ y(y-4) -8(y-4) } =y{ (y-8)(y-4) } =y(y-8)(y-4)

OpenStudy (anonymous):

thank you i have 5 more like that mind helping me out? @anikhalder and @jmprz_793

OpenStudy (anikhalder):

will try my best..need to show the working or just the answer?

OpenStudy (anonymous):

4c^2+4c-15 and 4x^2-2x^2-12x answer is fine thank you @anikhalder

OpenStudy (anikhalder):

first one is (2c-3)(2c+5) but in the second question is it 4x^2 or 4x^3? @JazminZamarripa

OpenStudy (anonymous):

sorry its 4x^3-2x^2-12x @anikhalder

OpenStudy (anonymous):

also x^3-512 its factoring

OpenStudy (anikhalder):

2nd answer is 2x(2x+3)(x-2) @JazminZamarripa

OpenStudy (anonymous):

x^3-512= (x-8)(x^2+8x+64

OpenStudy (anonymous):

)

OpenStudy (anikhalder):

512=8^3 so exp: x^3-8^3 =(x-8)(x^2 +8x + 64) (used the formula \[a^{3}-b^{3}=(a-b)(a ^{2} +ab + b ^{2})\]

OpenStudy (anonymous):

Thank you very very much can you help me on some more please=) @anikhalder

OpenStudy (anonymous):

i guess you don't want my help :( @JazminZamarripa haha

OpenStudy (anonymous):

oh sorry aha i thought you were offline=P want to help me? @jmprz_793

OpenStudy (anonymous):

Solve... How would you solve this? 2x^2+4x=6

OpenStudy (anonymous):

I have to find what x equals

OpenStudy (anonymous):

2x(x+2)=6 ;p x=-3 x=1

OpenStudy (anikhalder):

Thumbs up @jmprz_793

OpenStudy (anonymous):

aha=) thanks \[\frac{ x ^{2} }{ x ^{2}+6 }\times \frac{ x+5 }{ x }\]

OpenStudy (anonymous):

and \[\frac{ 5x ^{2}-25x }{ 3x-1 }\div \frac{ x-5 }{ 3x-1 }\]

OpenStudy (anonymous):

@jmprz_793 and @anikhalder

OpenStudy (anikhalder):

2nd answer is 5x

OpenStudy (anikhalder):

x(x+5)/(x^2+6) first answer Not sure about this

OpenStudy (anonymous):

\[\frac{ 9a ^{2} }{ 3a+7b }-\frac{ 49b ^{2} }{ 3a+7b }\]

OpenStudy (anikhalder):

(3a-7b)/(3a+7b)

OpenStudy (anonymous):

the first one is (x^(2))/(x^(2)+6)+(5x)/(x^(2)+6)

OpenStudy (anonymous):

Solve for x... \[\frac{ 3 }{ 1-x }=\frac{ 4 }{ 1+x}\]

OpenStudy (anonymous):

x=1/7

OpenStudy (anonymous):

@jmprz_793 how would i write that can you show me through the a drawing?

OpenStudy (anonymous):



OpenStudy (anikhalder):

3(1+x)=4(1-x) =>3+3x=4-4x =>7x=1 =>x=1/7

OpenStudy (anonymous):

Find domain of the function.... \[f(y)=\frac{ y+5 }{ y ^{2}-5y }\]

OpenStudy (anonymous):

Sorry i ment this one (x^(2))/(x^(2)+6)+(5x)/(x^(2)+6)

OpenStudy (anonymous):

domain: